A meter stick is projected into space at so a speed that it's length appears contracted to only 50cm. How fast is it going in m/s(2.6×10^8m/s)?


1
Expert's answer
2021-07-14T10:15:22-0400

Solution.

"l_0=1m; l=0.5m;"

"l=l_0\sqrt{1-\dfrac{v^2}{c^2}};"

"l^2=l_0^2(1-\dfrac{v^2}{c^2})\implies v=c\sqrt{1-\dfrac{l^2}{l_0^2}};"

"v=3\sdot10^8\sqrt{1-0.25}=2.6\sdot10^8m/s;"

Answer: "v=2.6\sdot10^8m/s."